OpenStudy (anonymous):

need help!?!?!?!?! will give medal Choose the equation below that represents the line passing through the point (1, −4) with a slope of one half. y − 4 = 1/2(x − 1) y + 4 = 2(x − 1) y + 4 = 1/2(x − 1) y − 4 = 2(x + 1)

OpenStudy (anonymous):

do you know your forms? Point-Slope, Standard and Slope-Intercept?

OpenStudy (anonymous):

yes but not sure how to apply them

OpenStudy (anonymous):

look at the info you are given... what info do you have?

OpenStudy (anonymous):

also, look at the answers... what form are the answers given in?

OpenStudy (anonymous):

the point (1,-4) and slope 1/2

OpenStudy (anonymous):

so you have a point and a slope... what form could you easliy put these into to get an equation for the line?

OpenStudy (anonymous):

point-slope?

OpenStudy (anonymous):

very good! please do so!

OpenStudy (anonymous):

y-y=m(x-x) so... y-(-4)=1/2(x-1)

OpenStudy (anonymous):

simplify a bit...

OpenStudy (anonymous):

y+4=1/2x-1/2

OpenStudy (anonymous):

too much... don't simplify the right side.

OpenStudy (anonymous):

ok,is the answer c then

OpenStudy (anonymous):

are you asking or telling?

OpenStudy (anonymous):

just making sure, I think the answer is c though

OpenStudy (anonymous):

well, be confident! justify and then testify!

OpenStudy (anonymous):

lol THE ANSWER IS C !!!!! thanks

OpenStudy (anonymous):

\(y-y_1=m\left( x-x_1\right)\) YES!!! AWESOME!!!